What is the wavelength of light that can cause a transition of an electron in the hydrogen atom from the orbit with n = 2 to n = 9?
Answer #1

The energy of the hydrogen atom can be written as, En = -13.6/n2 eV
Change in energy, E = E9 - E2
= -13.6 [1/92 - 1/22]
= 3.23 eV
= 3.23 * 1.60 * 10-16 J
= 5.17 * 10-16 J

E = hc/
= hc/(E)
= [6.63 * 10-34 * 3.00 * 108] / [5.17 * 10-16]
= 3.85 * 10-10 m